OpenStudy (anonymous):

integral of z^2 / (z-1)3

OpenStudy (ash2326):

\[\int \frac{z^2}{(z-1)^3} dz\] Substitute z=u+1 so \[dz=du\] We have now \[\int \frac{(u+1)^2}{(u)^3} dz\] Let's expand the square term \[\int \frac{(u^2+1+2u)}{(u)^3} dz\] we get \[\int \frac{1}{u}+ \frac{1}{u^3}+\frac{2}{u^2} du\] Can you do it now?

OpenStudy (anonymous):

but there is another way which i know and that's in which we do something like in numerator we place constants like A B C......

OpenStudy (ash2326):

That's the method of Partial fractions. Did you try it?

OpenStudy (anonymous):

the one which you have mentioned i dont used to solve partial fraction question...i used to take that method which i m telling you

OpenStudy (anonymous):

z^2= A/(z-1)^2 + B(z-1)

OpenStudy (anonymous):

this method will work, but lord is it a pain

OpenStudy (anonymous):

you will need \[\frac{z^2}{(z-1)^3}=\frac{a}{z-1}+\frac{b}{(z-1)^2}+\frac{c}{(z-1)^3}\] to do this

OpenStudy (anonymous):

\[\frac{z^2}{(z-1)^3}=\frac{1}{z-1}+\frac{2}{(z-1)^2}+\frac{1}{(z-1)^3} \]